# Jie Wu1∗  Haoling Li1∗  Xin Zhang2∗†‡  Jiani Guo3  Jane Luo2  Steven Liu2 Yangyu Huang2Ruihang Chu1‡Scarlett Li2Yujiu Yang1

### v1

@@ Abstract
Competitive programming presents great challenges for Code LLMs due to its intensive reasoning demands and high logical complexity. However, current Code LLMs still rely heavily on real-world data, which limits their scalability. In this paper, we explore a fully synthetic approach: training Code LLMs with entirely generated tasks, solutions, and test cases, to empower code reasoning models without relying on real-world data. To support this, we leverage feature-based synthesis to propose a novel data synthesis pipeline called SynthSmith. SynthSmith shows strong potential in producing diverse and challenging tasks, along with verified solutions and tests, supporting both supervised fine-tuning and reinforcement learning. Based on the proposed synthetic SFT and RL datasets, we introduce the X-Coder model series, which achieves a notable pass rate of 62.9 avg@8 on LiveCodeBench v5 and 55.8 on v6, outperforming DeepCoder-14B-Preview and AReal-boba²-14B despite having only 7B parameters. In-depth analysis reveals that scaling laws hold on our synthetic dataset, and we explore which dimensions are more effective to scale. We further provide insights into code-centric reinforcement learning and highlight the key factors that shape performance through detailed ablations and analysis. Our findings demonstrate that scaling high-quality synthetic data and adopting staged training can greatly advance code reasoning, while mitigating reliance on real-world coding data.
Figure 1: Left: SynthSmith generates high-quality synthetic tasks, solutions, and test cases to support both SFT and RL training. Right: Avg@8 results on LiveCodeBench. X-Coder achieves significant performance gains on competitive programming using fully synthetic data.
@@ 1 Introduction
As code language models advance, reasoning-focused models such as OpenAI-o1-ioi (OpenAI et al., 2025) have reached expert-level performance in programming. Classic benchmarks including HumanEval (Chen et al., 2021; Liu et al., 2023a) and MBPP (Austin et al., 2021) have been largely solved, whereas tasks from LiveCodeBench (Jain et al., 2024) and Codeforces continue to demand deeper reasoning and more complex algorithmic problem solving.
Recently, DeepSeek-R1 (Guo et al., 2025) has opened two opportunities for further boosting the reasoning capabilities of Code LLMs. The first is supervised fine-tuning (SFT) (Ouyang et al., 2022) on long Chain-of-Thought (CoT) demonstrations to distill reasoning patterns into student models (Hugging Face, 2025; Labs, 2025; Liu et al., 2025a). The second is reinforcement learning (RL) (Schulman et al., 2017) with GRPO (Shao et al., 2024b) and related algorithms to refine reasoning foundation models (Luo et al., 2025; Fu et al., 2025; He et al., 2025).
Both pathways have proven effective but face a common bottleneck: progress on competitive programming remains constrained by the scarcity of datasets. Widely used collections such as APPS (Hendrycks et al., 2021), CodeContests (Li et al., 2022), and TACO (Li et al., 2023) are heavily reused during post-training. They remain too modest in scale to support continued benefits and still lack the level of sufficiently challenging, diverse, and scalable. Meanwhile, collecting new real-world data tailored for competitive programming is also challenging. Although recent work has synthesized rewritten or evolutionary variants (Luo et al., 2024; Liu et al., 2025a; Xu et al., 2025) from existing resources, their diversity and complexity remain tightly bounded by the seed tasks.
To address this gap, we explore a fully synthetic approach: training Code LLMs with fully generated tasks, solutions, and test cases. Building on this insight, we present SynthSmith, a novel coding data synthesis pipeline tailored for competitive programming. To enable the synthesis of diverse and challenging competitive programming tasks, SynthSmith extends feature-based methods (Wang et al., 2025) with competition-oriented feature extraction, dedicated feature integration, and multi-style task construction. SynthSmith further supports the development of high-quality solutions and tool-based test case generation, both of which are cross-validated through the proposed dual-verification strategy. Thereby, SynthSmith demonstrates strong potential in producing scalable and challenging tasks, together with verified solutions and tests, offering support for both SFT and subsequent RL. Starting from a base model (e.g., Qwen3-8B-Base) or a non-reasoning model (e.g., Qwen2.5-Coder-7B-Instruct), we present the X-Coder series, which achieves significant performance gains on challenging LiveCodeBench v5 and v6 without relying on any real-world data, as shown in Figure 1. Beyond this, built upon verl (Sheng et al., 2025), we present an RL infrastructure featuring automated high-concurrency code validation, leveraging the CPUs of all distributed machines to support efficient and large-scale code execution.
Our in-depth analysis examines (i) whether synthetic SFT data scale effectively and which dimensions scale more favorably; (ii) the role of code-centric reinforcement learning, including the “good-gets-better” principle and RL’s resilience to noisy supervision; (iii) the factors that shape performance (long- vs. short-CoT, effects of solution verification, task style, and data-selection strategies); and (iv) the bottlenecks that limit code reasoning, together with the chained relationship among task difficulty, reasoning length, and pass rate. We further conduct case studies to uncover cognitive behaviors that emerge after SFT and RL, including reward hacking and undesirable patterns.
We make the following contributions:
(1) We explore a fully synthetic approach and propose a novel data synthesis pipeline tailored for competitive programming, producing high-quality datasets for both SFT and RL stages.
(2) We train both base and non-reasoning LLMs under an SFT-then-RL paradigm to develop the X-Coder model series, which achieves significant performance gains on LiveCodeBench v5 (avg@8: 62.9) and v6 (avg@8: 55.8), along with extensive analyses and ablations.
2 SynthSmith: Synthesis of Competition-Level Coding Data
Figure 2: Framework of SynthSmith. SynthSmith first extracts and evolves competitive programming related features from small-scale code instruction data and merges them into tree structures. It then samples subtrees from the feature tree, selects a compatible feature set, and formulates a scenario that naturally integrates these consistent features. Novel tasks are generated based on a proposed scenario according to specific styles. Advanced reasoning models are used to synthesize solutions and tests for the generated tasks, which are further cross-verified using the proposed dual-verification strategy to yield reliable test outputs and the top solution.
@@ 2 SynthSmith: Synthesis of Competition-Level Coding Data
We introduce SynthSmith, a fully synthetic framework for constructing competitive programming tasks that support both the SFT and RL stages. Figure 2 illustrates the SynthSmith pipeline, which consists of four key steps: (i) generating novel and challenging problems (with the capacity for easy scaling in quantity); (ii) constructing diverse and comprehensive input test cases for each problem (including boundary and stress tests); (iii) producing high-quality candidate solutions; and (iv) employing a dual-verification strategy that cross-checks solutions with test cases to yield more accurate test outputs and more reliable solutions.
@@
(i) Task Generation. Inspired by EpiCoder (Wang et al., 2025), which generates novel programming tasks through a feature-based framework by combining sampled features into problem scenarios, we extend this approach with three key improvements to synthesize diverse and complex tasks tailored for competitive programming. First, instead of relying on broad definitions of features, we explicitly extract and evolve competition-related features from 10k code snippets in the TACO dataset (Li et al., 2023) using GPT-4o-0513 (detailed in §B.1). Second, formulating competitive scenarios from a rich feature tree is non-trivial, as LLMs often oversimplify complex prompts into trivial cases, thereby reducing both diversity and difficulty. To address this, we adopt a two-stage process that separates feature selection from scenario formulation: first, selecting mutually consistent features for meaningful composition; and second, formulating hint-free tasks that demand genuine reasoning. We further incorporate one-shot prompting to improve task understanding and instruction-following. Third, we adapt the synthesis method to support multi-style task generation, covering Codeforces [^1] https://codeforces.com/-style tasks (standard input/output with imaginable narrative contexts), LeetCode-style [^2] https://leetcode.com/ tasks (starter code with predefined function signatures), and AtCoder [^3] https://atcoder.jp/-style tasks (concise specifications with minimal explanations), thereby enhancing task diversity. Examples of the task generation process are provided in §B.2, together with difficulty estimates on generated tasks in §B.3.
@@
(ii) Test Input Generation. Obtaining sufficient and accurate test cases is a formidable challenge. Problems from competitive programming platforms often do not provide test cases, or only provide a limited number, due to platform constraints. This results in insufficient quantity, difficulty, and coverage of test cases during RL training. To address the inherent scarcity of test cases in synthesized data, we investigate two complementary methods for generating the input component of the test case. The Prompting-based method instructs the LLM to interpret the problem’s input constraints and directly generate multiple test inputs, covering both standard cases and edge-case instances. The Tool-based method leverages CYaRon [^4] https://github.com/luogu-dev/cyaron, a dedicated test case generation library, enabling the LLM to construct test inputs by invoking functions documented within the library after understanding the problem. For each task, we generate a set of $n$ test case inputs $x_{1},x_{2},\ldots,x_{n}]$. Detailed description of test input generation is provided in §D, and a comparative analysis is presented in Sec 6.
@@
(iii) Candidate Solutions Generation. For each task, we generate multiple candidate solutions using advanced open-source reasoning LLMs, obtaining $m$ answers $A^{1},A^{2},\ldots,A^{m}]$. We verify that each candidate solution includes a complete reasoning process and a Python code implementation, and we ensure the absence of syntax errors through static analysis methods based on Abstract Syntax Tree (AST). Filtering criteria are provided in §C.1.
(iv) Dual-Verification of Solutions and Test Cases.
@@
To ensure the robustness and reliability of both the generated solutions and the constructed test cases, we adopt a dual-verification strategy. Step 1 of this strategy extends the principle of self-consistency (Wang et al., 2023) by applying majority voting across candidate solutions from multiple LLMs, which mitigates model-specific biases and enhances generalization, thereby yielding a reliable test output for each input. Step 2 then identifies the top-performing candidate solution by incorporating test case difficulty weighting alongside a hold-out validation set.
@@
where $\mathbb{I}(\cdot)$ is the indicator function. This yields a candidate test set $\mathcal{T}_{candidate}=\{(x_{1},\hat{y}_{1}),\ldots,(x_{n},\hat{y}_{n})\}$. Crucially, we posit that not all test cases are of equal importance; boundary or edge cases are critical for robust evaluation. We therefore introduce a weighting function $w(x_{i})\to w_{i}$ that assigns a higher score to more challenging test cases. The weight $w_{i}$ is determined by a set of heuristics based on input characteristics, such as character or token count, structural complexity, or semantic novelty, which serve as proxies for difficulty.
@@
$A^{\prime}_{golden}=\underset{A^{j}\in\{A^{1},\ldots,A^{m}\}}{\operatorname{argmax}}\sum_{(x_{i},\hat{y}_{i})\in\mathcal{T}_{golden}}w_{i}\cdot\mathbb{I}(A^{j}(x_{i})=\hat{y}_{i})\quad.$ (2)
The final confirmation of $A_{golden}$ is contingent upon its performance on the unseen hold-out set $\mathcal{T}_{val}$. We verify that $A^{\prime}_{golden}$ also achieves the highest (or a competitively high) unweighted accuracy on $\mathcal{T}_{val}$ relative to other candidates. This additional validation step ensures that the selected solution is not merely overfitted to the specifics of the weighted test cases but demonstrates superior, generalizable correctness. The detailed algorithm is provided in §E.
Finally, we obtain $A_{golden}$ and ${T}_{golden}$ for each task $q$. The pair $q,A_{golden}$}] is used to compute the SFT loss, and $q,T_{golden}$}] are used for RL via the GRPO algorithm.
Discussion. Compared to rewriting-based data synthesis methods (Luo et al., 2024; Liu et al., 2025a), SynthSmith reduces reliance on seed tasks by formulating novel tasks from evolved competitive features. Compared with EpiCoder, it generates more challenging tasks and selects high-quality solutions via a dual-verification strategy, yielding a 21% absolute performance gain on LiveCodeBench v5 (Figure 5(c)). Moreover, SynthSmith extends data synthesis to the RL stage, showing that synthetic RL data can further improve performance beyond the SFT model as shown in Table 1.

### v2

# Synthetic Tasks, Solutions, and Tests

@@ Abstract
Competitive programming poses a significant challenge for Code LLMs. While recent models have shown promise, they heavily rely on finite real-world data, raising concerns about scalability and contamination. In this paper, we investigate a critical question: Can we elevate models to expert-level reasoning performance using fully synthetic data? In response, we first observe that off-the-shelf synthesis methods yield suboptimal results in this domain. To address this, we systematically investigate the key factors governing synthetic data quality. Leveraging these findings, we significantly advance the feature-based synthesis paradigm via domain-specific evolution and a dual-verification strategy, promoting task solvability, solution correctness, and test accuracy. Using this high-quality synthetic data, we train the X-Coder model series under an SFT-then-RL paradigm. X-Coder-7B shows significant performance gains on the challenging LiveCodeBench v5 (62.9% avg@8) and v6 (55.8% avg@8), outperforming larger models trained on real-world data. Extensive analysis distills valuable insights into synthetic data scaling, the necessity of domain-adapted feature evolution, and code-centric reinforcement.
Figure 1: Left: To address persistent data scarcity in competitive programming, we build a scalable alternative by training Code LLMs with fully synthetic data. Right: Avg@8 results on LiveCodeBench v5. X-Coder achieves significant performance gains on competitive programming using fully synthetic data.
@@ 1 Introduction
As Code LLMs advance, they have demonstrated remarkable proficiency in basic programming tasks. Benchmarks such as HumanEval (Chen et al., 2021; Liu et al., 2023a) and MBPP (Austin et al., 2021) are largely considered solved. However, the frontier of code intelligence has shifted towards competitive programming. This domain is exemplified by platforms like Codeforces and benchmarks like LiveCodeBench (Jain et al., 2024), requiring deep algorithmic reasoning and complex problem-solving capabilities.
To bridge this gap, recent success in reasoning models has been driven by the utilization of extensive reasoning data, both for supervised fine-tuning (SFT) and reinforcement learning (RL) (Guo et al., 2025; Hugging Face, 2025; Luo et al., 2025). However, this data-driven scaling faces a critical bottleneck: the scarcity of high-quality training data. The finite pool of real-world competitive programming tasks is insufficient to sustain the scaling laws of reasoning models. Furthermore, existing collections (Hendrycks et al., 2021; Li et al., 2022) are heavily exhausted, raising concerns about data contamination and overfitting. This dilemma poses a critical question: Can we bypass the reliance on real-world data and elevate models to expert-level reasoning performance using fully synthetic data?
Answering this question is non-trivial. Generating high-quality synthetic data for competitive programming is significantly harder than for general-purpose programming, facing unique challenges across three dimensions:
(i) Tasks must be both solvable and challenging. Synthesizing competitive-level tasks often yields ill-defined or trivial instances, failing to provide the rigorous complexity needed to train expert models. (ii) High-quality SFT demands logically sound solutions. Validating these without ground-truth test cases is inherently risky, potentially polluting the training set with incorrect reasoning paths. (iii) RL necessitates reliable reward signals. Weak or insufficient synthetic tests fail to catch incorrect solutions, providing noisy reward signals that can mislead the reinforcement learning process.
Given these challenges, off-the-shelf synthesis frameworks (Wang et al., 2025; Wei et al., 2024) fail to generalize to competitive programming and yield suboptimal results. In this work, we validate the hypothesis that fully synthetic data can sufficiently drive expert-level performance in code reasoning. To achieve this, we systematically investigate the key factors governing the effectiveness of synthetic code reasoning data. Leveraging these findings, we significantly advance the synthesis paradigm via domain-specific evolution and a dual-verification strategy, establishing a high-quality synthetic foundation for training.
To validate the feasibility of fully synthetic training, we train the X-Coder model series using only the synthesized data under a standard SFT-then-RL paradigm. As shown in Figure 1, X-Coder achieves significant performance gains on the challenging LiveCodeBench v5 and v6.
Through extensive evaluation, we further distill key insights into what makes synthetic data effective, shedding light on synthetic data scaling, the necessity of domain-specific evolution, and the role of code-centric reinforcement learning.
Our contributions are as follows:
(1) We promote synthetic data quality via domain-specific adaptation and dual-verification, empirically demonstrating that fully synthetic data is sufficient to achieve expert-level performance in competitive programming, challenging the conventional reliance on real-world datasets.
(2) Using the constructed synthetic data, we develop the X-Coder model series under an SFT-then-RL paradigm. X-Coder-7B achieves significant performance gains on LiveCodeBench v5 and v6, particularly on medium and hard splits. Related resources will be made publicly available.
(3) We offer insights into synthetic data scaling, highlighting the necessity of domain-adapted feature evolution and dual-verification. Furthermore, we shed light on reasoning types, code-centric reinforcement learning, and emergent fine-grained failure modes.
2 Related Work
Data Synthesis for Code Generation. The research community has long recognized the scarcity of high-quality coding tasks. To address this, Wizard-Coder (Luo et al., 2024) extends Evol-Instruct (Xu et al., 2024) by evolving basic code-instruction data into augmented variants. rStar-Coder (Liu et al., 2025a) further adapts this augmentation strategy to the competitive programming domain. CodeEvo (Sun et al., 2025) introduces a coder–reviewer interaction framework to collaboratively synthesize high-quality instruction–code pairs. SelfCodeAlign (Wei et al., 2024) pioneered concept composition, generating novel tasks by systematically recombining fundamental concepts extracted from seed problems. EpiCoder (Wang et al., 2025) follows this direction by sampling sub-features from a large and expressive feature tree to formulate novel problems, further improving task complexity and diversity.
However, applying these general-purpose techniques to competitive programming yields suboptimal results. To bridge this gap, we systematically investigate the key factors governing data effectiveness and develop a synthesis strategy tailored for code reasoning. Specifically, we advance feature-based synthesis via domain-specific evolution to generate complex tasks and employ a dual-verification strategy to guarantee solution correctness. This approach establishes a scalable, high-quality, fully synthetic foundation capable of driving expert-level performance.
Post-training for Competitive Programming. Current training paradigms generally fall into three categories: (i) Purely supervised fine-tuning (SFT) on real-world tasks or their variants (Labs, 2025; Guha et al., 2025; Liu et al., 2025a); (ii) Purely reinforcement learning (RL) using GRPO-like algorithms (Shao et al., 2024b; He et al., 2025; Luo et al., 2025; Fu et al., 2025); and (iii) A hybrid SFT-then-RL approach, which typically mixes coding and mathematical data to mitigate the scarcity of high-quality code problems (Liu et al., 2025b; Xiaomi et al., 2025; Su et al., 2025). Crucially, all these approaches rely heavily on finite real-world data sources. The hybrid methods, in particular, often depend on mathematical reasoning transfer due to the lack of sufficient competitive programming data. In contrast, we demonstrate that fully synthetic data can independently sustain the full SFT-then-RL cycle for code, achieving significant performance without relying on real-world data or cross-domain mixtures.
3 Synthesis of Competition-Level Coding Data
Figure 2: Overview of the data synthesis framework. The framework begins with feature extraction and evolution to construct a domain-specific feature tree. From this, we generate multi-style competitive tasks by formulating scenarios from compatible feature sets. Finally, we synthesize solutions and test inputs, employing a dual-verification strategy to ensure solution correctness and test accuracy, yielding a high-quality data foundation for subsequent SFT and RL.
@@ 3 Synthesis of Competition-Level Coding Data
We adapt the feature-based synthesis framework (Wang et al., 2025) to the competitive programming domain, establishing a comprehensive methodology to construct tasks that support both SFT and RL stages. As illustrated in Figure 2, this process addresses the unique challenges of synthesis through four key steps: (i) generating novel and challenging problems (with the capacity for easy scaling in quantity); (ii) constructing diverse and comprehensive test inputs for each problem (including boundary and stress tests); (iii) producing high-quality candidate solutions; and (iv) employing a dual-verification strategy that cross-checks solutions with test cases to yield more accurate test outputs and more reliable solutions.
@@
(i) Task Generation. While EpiCoder (Wang et al., 2025) introduced feature-based synthesis for general coding tasks, we extend this approach with three key improvements specifically tailored for the complexity of competitive programming. First, instead of relying on broad definitions of features, we explicitly extract and evolve competition-related features from 10k code snippets in the TACO dataset (Li et al., 2023) using GPT-4o-0513 (detailed in Appendix B.1). Second, formulating competitive scenarios from a rich feature tree is non-trivial, as LLMs often oversimplify complex prompts into trivial cases, thereby reducing both diversity and difficulty. To address this, we adopt a two-stage process that separates feature selection from scenario formulation: first, selecting mutually consistent features for meaningful composition; and second, formulating hint-free tasks that demand genuine reasoning. This two-stage approach significantly outperforms single-step generation (Appendix Table 9). We further incorporate one-shot prompting to improve task understanding and instruction-following. Third, we adapt the synthesis method to support multi-style task generation, covering Codeforces [^1] https://codeforces.com/-style tasks (standard input/output with narrative-rich contexts), LeetCode-style [^2] https://leetcode.com/ tasks (starter code with predefined function signatures), and AtCoder [^3] https://atcoder.jp/-style tasks (concise specifications with minimal explanations), thereby enhancing task diversity. Examples of the task generation process are provided in Appendix B.2, together with difficulty estimates on generated tasks in Appendix B.3 and style comparisons in Appendix B.2.3.
@@
(ii) Test Input Generation. Obtaining sufficient and accurate test cases is a formidable challenge. Problems from competitive programming platforms often do not provide test cases, or only provide a limited number, due to platform constraints. This results in insufficient quantity, difficulty, and coverage of test cases during RL training. To address the inherent scarcity of test cases in synthesized data, we investigate two complementary methods for generating the input component of the test case. The Prompting-based method instructs the LLM to interpret the problem’s input constraints and directly generate multiple test inputs, covering both standard cases and edge-case instances. The Tool-based method leverages CYaRon [^4] https://github.com/luogu-dev/cyaron, a dedicated test case generation library, enabling the LLM to construct test inputs by invoking functions documented within the library after understanding the problem. For each task, we generate a set of $n$ test case inputs $x_{1},x_{2},\ldots,x_{n}]$. Detailed description of test input generation is provided in Appendix D, and a comparative analysis is presented in Appendix D.2.
@@
(iii) Candidate Solutions Generation. For each task, we generate multiple candidate solutions using advanced open-source reasoning LLMs, obtaining $m$ answers $A^{1},A^{2},\ldots,A^{m}]$. We verify that each candidate solution includes a complete reasoning process and a Python code implementation, and we ensure the absence of syntax errors through static analysis methods based on Abstract Syntax Tree (AST). Filtering criteria are provided in Appendix C.1.
@@
(iv) Dual-Verification of Solutions and Test Cases. We adopt a dual-verification strategy. Step 1 of this strategy extends the principle of self-consistency (Wang et al., 2023) by applying majority voting across candidate solutions from multiple LLMs, which mitigates model-specific biases and enhances generalization, thereby yielding a reliable test output for each input. Step 2 then identifies the top-performing candidate solution by incorporating test case difficulty weighting alongside a hold-out validation set.
@@
where $\mathbb{I}(\cdot)$ is the indicator function. Empirical evaluation on the TACO dataset demonstrates that this voting strategy achieves 94.7% labeling accuracy with 8 sampled solutions (see analysis in Appendix E.3 and Table 14). This yields a candidate test set $\mathcal{T}_{candidate}=\{(x_{1},\hat{y}_{1}),\ldots,(x_{n},\hat{y}_{n})\}$. Crucially, we posit that not all test cases are of equal importance; boundary or edge cases are critical for robust evaluation. We therefore introduce a weighting function $w(x_{i})\to w_{i}$ that assigns a higher score to more challenging test cases. The weight $w_{i}$ is determined by combining semantic-based heuristics (e.g., assigning higher weights to boundary and stress tests explicitly requested in the prompt) and size-based metrics (e.g., input file size as a proxy for computational complexity). Detailed weighting criteria are provided in Appendix E.2.
@@
$A^{\prime}_{golden}=\underset{A^{j}\in\{A^{1},\ldots,A^{m}\}}{\operatorname{argmax}}\sum_{(x_{i},\hat{y}_{i})\in\mathcal{T}_{golden}}w_{i}\cdot\mathbb{I}(A^{j}(x_{i})=\hat{y}_{i})$
The final confirmation of $A_{golden}$ is contingent upon its performance on the unseen hold-out set $\mathcal{T}_{val}$. We verify that $A^{\prime}_{golden}$ also achieves the highest (or a competitively high) unweighted accuracy on $\mathcal{T}_{val}$ relative to other candidates. This additional validation step ensures that the selected solution is not merely overfitted to the specifics of the weighted test cases but demonstrates superior, generalizable correctness. The detailed algorithm is provided in Appendix E.1.
We ensure task solvability by using GPT-5 (high reasoning effort) as a proxy solver, filtering out approximately 36.9% of tasks where it achieves a zero pass rate on the voted test cases. We posit that if such a capable model fails completely, the task likely suffers from ambiguity or underspecification. The pass rate distribution of the remaining tasks is shown in Table 18 (Appendix E.5).
Ultimately, this process yields a verified solution $A_{golden}$ and a comprehensive test suite $\mathcal{T}_{golden}$ for every generated task $q$. We leverage these high-quality synthetic assets to drive post-training for Code LLMs: $(q,A_{golden})$ pairs for supervised fine-tuning, and $(q,\mathcal{T}_{golden})$ for reinforcement learning via GRPO algorithm (Shao et al., 2024b).